# LASER: Layer-wise Scale Alignment for Training-Free Streaming 4D Reconstruction

### v1

@@ 4.1.2 Baselines
Offline feed-forward models. We include DUSt3R [dust3r], Fast3R [Yang_2025_Fast3R], VGGT [wang2025vggt], and $\pi^{3}$ [wang2025pi], which process static image batches without temporal constraints.
@@
Classical SLAM systems. For camera pose evaluation on KITTI Odometry, we compare to SLAM methods including ORB-SLAM2 [murORB2], LDSO [gao2018ldso], DROID-VO, DROID-SLAM [teed2021droid], DPV-SLAM, and DPV-SLAM++ [lipson2024deep].
Training-free concurrent work. To further demonstrate the strength of our training-free design and ensure that performance is not dominated by a strong backbone, we also evaluate against VGGT-Long [deng2025vggtlongchunkitloop], a concurrent training-free streaming framework built on VGGT [wang2025vggt]. For a fair comparison, we re-implement its pipeline on the $\pi^{3}$ [wang2025pi] backbone, denoted as $\pi^{3}$-Long, enabling a one-to-one comparison under identical base models.
@@ 4.1.3 Implementation Details
We instantiate LASER using either VGGT [wang2025vggt] or $\pi^{3}$ [wang2025pi] as the offline 4D reconstruction backbone. On the kilometer-scale KITTI Odometry, we incorporate loop closure following the VGGT-Long [deng2025vggtlongchunkitloop] configuration for fair comparison. More details are in the supplemental material.
@@ 4.3 Camera Pose Estimation
On large-scale outdoor sequences (Tab. 3), LASER using $\pi^{3}$ as backbone achieves the second-lowest mean ATE among all methods on both Avg. and Avg. metrics. It attains accuracy comparable to or better than well-designed SLAM systems such as ORB-SLAM2 [murORB2] and DROID-SLAM [teed2021droid], which yield only sparse reconstructions and may require camera calibration. Meanwhile, dense offline models like VGGT [wang2025vggt] and $\pi^{3}$ [wang2025pi] fail to process long sequences due to memory limits, and streaming variants such as CUT3R [wang2025cut3r] and MASt3R-SLAM [mast3r_slam] either run out of memory or lose tracking. In contrast, LASER remains stable across all eleven sequences, producing globally consistent trajectories. LASER also outperforms training-free streaming concurrent work VGGT-Long [deng2025vggtlongchunkitloop] and its variant $\pi^{3}$-Long by a $2121\%$ reduction in avg ATE. Notably, although $\pi^{3}$-Long shares the same backbone with ours, it performs worse, indicating that the improvement is from our streaming algorithm design rather than backbone capacity.
@@ 4.5 Qualitative Results
Figure 6: Qualitative comparisons on DAVIS [davis] and Hike [meuleman2023localrf] (top to bottom). For each sequence, we show the reconstructed global point cloud with the estimated camera trajectory overlaid. Please zoom in for camera trajectory details.
@@ 4.6 Efficiency Analysis
LASER also demonstrates strong efficiency, as shown in Fig. 5. All experiments are performed on a RTX A6000 GPU. Compared to streaming feed-forward baselines, our method achieves the highest runtime speed ($\sim$14.2 FPS) with only 6 GB of peak memory usage when using $\pi^{3}$ [wang2025pi] as our offline model, while maintaining superior performance in video depth and camera pose estimation. When using VGGT [wang2025vggt], we also have a competitive inference speed of $\sim$10.9 FPS and 10 GB peak memory usage.
@@ 4.7 Ablation Studies
We evaluate key components of our pipeline through a series of ablations, using $\pi^{3}$ [wang2025pi] as the backbone.
Layer-wise Scale Alignment (Sec. 3.2). Tab. 5 investigates the components of the LSA module on video depth estimation, as LSA does not affect camera pose estimation. Disabling LSA leads to clear drops in depth accuracy. We also try substituting the segmentation algorithm [felzenszwalb2004efficient] with SAM 2 [ravi2024sam2]. Despite trading speed for better segmentation, SAM 2 does not improve accuracy. Finally, disabling propagation through $\mathcal{E}_{\text{intra}}$ ignores temporal relationships and prevents scale updates in non-overlapping frames, which harms global consistency across long sequences.
@@ 7 More Details about Submap Registration in $\mathrm{Sim}(3)$ Space
To estimate the $\mathrm{Sim}(3)$ transform, we first estimate the global scale factor $s_{i}^{w}$ via a robust IRLS (Iteratively Reweighted Least Squares) optimization, enforcing a shared metric across two adjacent windows. Rotation and translation $(\mathbf{R}_{i}^{w},\mathbf{t}_{i}^{w})$ are then optimized via the Kabsch algorithm  [Kabsch] under that metric using the scaled camera anchors based on the estimated $s_{i}^{w}$.
@@
We then estimate the window-level rigid transform $(\mathbf{R}_{i}^{w},\mathbf{t}_{i}^{w})$ by minimizing the alignment error between the two anchor sets via the Kabsch algorithm [Kabsch]:
@@
Differences from existing approaches. Although VGGT-Long [deng2025vggtlongchunkitloop] also adopts a sliding-window strategy for streaming inputs, our method differs in how the registration $\mathrm{Sim}(3)$ is estimated from overlapping windows. VGGT-Long applies IRLS to jointly optimize a closed-form scale $s$ together with $\mathbf{R}$ and $\mathbf{t}$ computed via the Kabsch algorithm. In contrast, we first estimate the scale using point-cloud correspondences within matched camera coordinate systems, and then estimate $\mathbf{R}$ and $\mathbf{t}$ using the scaled inputs. This two-stage procedure yields more stable and robust registration.
@@
Table 7: Outdoor, Long-term Point Map Estimation on Waymo [waymo]. We report Accuracy (Acc, lower is better), Completeness (Comp, lower is better) and Chamfer Distance (Chamfer, lower is better). We show metrics for each segment ID; Avg. is the mean across segments.
@@
VGGT–Long [deng2025vggtlongchunkitloop] Acc $\downarrow$ 0.508 0.453 0.096 0.629 1.441 0.457 0.379 0.510 0.481 0.129
@@ 8 Implementation Details
We instantiate LASER using either VGGT [wang2025vggt] or $\pi^{3}$ [wang2025pi] as the offline 3D reconstruction backbone. For video depth estimation, small-scale camera pose estimation, and indoor multi-view point map estimation, we evaluate both variants. For large-scale camera pose estimation on KITTI Odometry [kitti] and outdoor point map estimation on Waymo [waymo], we use $\pi^{3}$ as the backbone for its stronger geometric prior. On the kilometer-scale KITTI Odometry, we additionally incorporate loop closure following the VGGT-Long [deng2025vggtlongchunkitloop] configuration for fairness.
@@
Fig. 9 and Tab. 8 compares alternative strategies for estimating the $\mathrm{SE}(3)$ transform between submaps. Replacing IRLS with a closed-form solver degrades both depth and pose accuracy, confirming the importance of robust scale estimation in this stage. Replacing scaled camera anchors with scaled point maps produces similar depth metrics but noticeably weaker camera trajectories.
@@ 11 Evaluation Details of Efficiency Benchmark
We report FPS and peak memory usage on the Sintel [sintel] benchmark for all methods on an A6000 GPU. The image resolution for DUSt3R-based [dust3r] methods is $512\times 288$ except Spann3R, which only supports $224\times 224$, and VGGT-based [wang2025vggt] methods are $518\times 294$.
@@ 12 Outdoor Multi-view Point Map Estimation
Tab. 7 reports long-term multi-view point map estimation results. LASER using $\pi^{3}$ as backbone achieves the best overall performance among training-free methods, substantially outperforming both VGGT-Long [deng2025vggtlongchunkitloop] and $\pi^{3}$-Long in Comp and Chamfer while maintaining comparable Acc.
For outdoor setting, we use the Waymo Open Dataset [waymo] on urban driving segments and report Acc, Comp, and Chamfer distance, following [deng2025vggtlongchunkitloop] (results in the supplementary material). To mitigate artifacts from sky and far-background regions, we uniformly filter out the lowest-confidence 40% of predicted points for all methods; these results serve as a comparative reference rather than a strict head-to-head benchmark.
@@
CUT3R [wang2025cut3r]
StreamVGGT [zhuo2025streaming]
VGGT-SLAM [vggtslam]
@@ 13 Future Directions
Although our method demonstrates strong performance, it has room for improvements. We show some failure cases in Fig. 11, and list two directions that interest us most:

### v2

@@ 4.1.2 Baselines
Offline feed-forward models. We include DUSt3R [63], Fast3R [68], VGGT [59], and $\pi^{3}$ [64], which process static image batches without temporal constraints.
@@
Classical SLAM systems. For camera pose evaluation on KITTI Odometry, we compare to SLAM methods including ORB-SLAM2 [38], LDSO [15], DROID-VO, DROID-SLAM [55], DPV-SLAM, and DPV-SLAM++ [31].
Training-free concurrent work. To further demonstrate the strength of our training-free design and ensure that performance is not dominated by a strong backbone, we also evaluate against VGGT-Long [9], a concurrent training-free streaming framework built on VGGT [59]. For a fair comparison, we re-implement its pipeline on the $\pi^{3}$ [64] backbone, denoted as $\pi^{3}$-Long, enabling a one-to-one comparison under identical base models.
@@ 4.1.3 Implementation Details
We instantiate LASER using either VGGT [59] or $\pi^{3}$ [64] as the offline 4D reconstruction backbone. On the kilometer-scale KITTI Odometry, we incorporate loop closure following the VGGT-Long [9] configuration for fair comparison. More details are in the supplemental material.
@@ 4.2 Video Depth Estimation
Tab. 1 shows the video depth estimation results. For streaming methods, a slight degradation on performance is reasonable compared with offline methods, since streaming methods operate with limited temporal context which leads to inconsistent depth predictions.
@@ 4.3 Camera Pose Estimation
On large-scale outdoor sequences (Tab. 3), LASER using $\pi^{3}$ as backbone achieves the second-lowest mean ATE among all methods on both Avg. and Avg. metrics. It attains accuracy comparable to or better than well-designed SLAM systems such as ORB-SLAM2 [38] and DROID-SLAM [55], which yield only sparse reconstructions and may require camera calibration. Meanwhile, dense offline models like VGGT [59] and $\pi^{3}$ [64] fail to process long sequences due to memory limits, and streaming variants such as CUT3R [62] and MASt3R-SLAM [39] either run out of memory or lose tracking. In contrast, LASER remains stable across all eleven sequences, producing globally consistent trajectories. LASER also outperforms training-free streaming concurrent work VGGT-Long [9] and its variant $\pi^{3}$-Long by a $2121\%$ reduction in avg ATE. Notably, although $\pi^{3}$-Long shares the same backbone with ours, it performs worse, indicating that the improvement is from our streaming algorithm design rather than backbone capacity.
@@ 4.5 Qualitative Results
Figure 6: Qualitative comparisons on DAVIS [45] and Hike [36] (top to bottom). For each sequence, we show the reconstructed global point cloud with the estimated camera trajectory overlaid. Please zoom in for camera trajectory details.
@@ 4.6 Efficiency Analysis
LASER also demonstrates strong efficiency, as shown in Fig. 5. All experiments are performed on a RTX A6000 GPU. Compared to streaming feed-forward baselines, our method achieves the highest runtime speed ($\sim$14.2 FPS) with only 6 GB of peak memory usage when using $\pi^{3}$ [64] as our offline model, while maintaining superior performance in video depth and camera pose estimation. When using VGGT [59], we also have a competitive inference speed of $\sim$10.9 FPS and 10 GB peak memory usage.
@@ 4.7 Ablation Studies
We evaluate key components of our pipeline through a series of ablations, using $\pi^{3}$ [64] as the backbone.
Layer-wise Scale Alignment (Sec. 3.2). Tab. 5 investigates the components of the LSA module on video depth estimation, as LSA does not affect camera pose estimation. Disabling LSA leads to clear drops in depth accuracy. We also try substituting the segmentation algorithm [11] with SAM 2 [47]. Despite trading speed for better segmentation, SAM 2 does not improve accuracy. Finally, disabling propagation through $\mathcal{E}_{\text{intra}}$ ignores temporal relationships and prevents scale updates in non-overlapping frames, which harms global consistency across long sequences.
@@ 7 More Details about Submap Registration in $\mathrm{Sim}(3)$ Space
To estimate the $\mathrm{Sim}(3)$ transform, we first estimate the global scale factor $s_{i}^{w}$ via a robust IRLS (Iteratively Reweighted Least Squares) optimization, enforcing a shared metric across two adjacent windows. Rotation and translation $(\mathbf{R}_{i}^{w},\mathbf{t}_{i}^{w})$ are then optimized via the Kabsch algorithm  [22] under that metric using the scaled camera anchors based on the estimated $s_{i}^{w}$.
@@
We then estimate the window-level rigid transform $(\mathbf{R}_{i}^{w},\mathbf{t}_{i}^{w})$ by minimizing the alignment error between the two anchor sets via the Kabsch algorithm [22]:
@@
Differences from existing approaches. Although VGGT-Long [9] also adopts a sliding-window strategy for streaming inputs, our method differs in how the registration $\mathrm{Sim}(3)$ is estimated from overlapping windows. VGGT-Long applies IRLS to jointly optimize a closed-form scale $s$ together with $\mathbf{R}$ and $\mathbf{t}$ computed via the Kabsch algorithm. In contrast, we first estimate the scale using point-cloud correspondences within matched camera coordinate systems, and then estimate $\mathbf{R}$ and $\mathbf{t}$ using the scaled inputs. This two-stage procedure yields more stable and robust registration.
@@
Table 7: Outdoor, Long-term Point Map Estimation on Waymo [54]. We report Accuracy (Acc, lower is better), Completeness (Comp, lower is better) and Chamfer Distance (Chamfer, lower is better). We show metrics for each segment ID; Avg. is the mean across segments.
@@
VGGT–Long [9] Acc $\downarrow$ 0.508 0.453 0.096 0.629 1.441 0.457 0.379 0.510 0.481 0.129
@@ 8 Implementation Details
We instantiate LASER using either VGGT [59] or $\pi^{3}$ [64] as the offline 3D reconstruction backbone. For video depth estimation, small-scale camera pose estimation, and indoor multi-view point map estimation, we evaluate both variants. For large-scale camera pose estimation on KITTI Odometry [16] and outdoor point map estimation on Waymo [54], we use $\pi^{3}$ as the backbone for its stronger geometric prior. On the kilometer-scale KITTI Odometry, we additionally incorporate loop closure following the VGGT-Long [9] configuration for fairness.
@@
Fig. 9 examines the effect of selecting different overlap sizes $O$, our method is robust to a wide range of settings, the default choice ($O=5$) maintains a good balance between accuracy and inference complexity. Fig. 10 and Tab. 8 compares alternative strategies for estimating the $\mathrm{SE}(3)$ transform between submaps. Replacing IRLS with a closed-form solver degrades both depth and pose accuracy, confirming the importance of robust scale estimation in this stage. Replacing scaled camera anchors with scaled point maps produces similar depth metrics but noticeably weaker camera trajectories.
@@ 11 Evaluation Details of Efficiency Benchmark
We report FPS and peak memory usage on the Sintel [2] benchmark for all methods on an A6000 GPU. The image resolution for DUSt3R-based [63] methods is $512\times 288$ except Spann3R, which only supports $224\times 224$, and VGGT-based [59] methods are $518\times 294$.
@@ 12 Outdoor Multi-view Point Map Estimation
Tab. 7 reports long-term multi-view point map estimation results. LASER using $\pi^{3}$ as backbone achieves the best overall performance among training-free methods, substantially outperforming both VGGT-Long [9] and $\pi^{3}$-Long in Comp and Chamfer while maintaining comparable Acc.
For outdoor setting, we use the Waymo Open Dataset [54] on urban driving segments and report Acc, Comp, and Chamfer distance, following [9] (results in the supplementary material). To mitigate artifacts from sky and far-background regions, we uniformly filter out the lowest-confidence 40% of predicted points for all methods; these results serve as a comparative reference rather than a strict head-to-head benchmark.
@@
CUT3R [62]
StreamVGGT [73]
VGGT-SLAM [35]
@@ 13 Future Directions
Although our method demonstrates strong performance, it has room for improvements. We show some failure cases in Fig. 12, and list three directions that interest us most:
@@
  • LSA Sensitivity to complex scenes and object occlusions. Our current depth layer graph construction used in LSA is based purely on predicted frame-wise depth maps, without persistent contexts of distinct objects, which results in multiple objects being assigned to the same depth layer within complex scenes, also unable to recover layer correspondences if an object is occluded entirely and later re-enters the scene. We seek to improve the robustness of our design and the temporal trackability of different objects by incorporating appearance information from raw images.